a boomerang is thrown with the following displacements: 4.5-m to the east, 7.5-m 30° north of west, and 3.0-m 10° south of west. what is the resulatant displacement?
1
Expert's answer
2020-06-02T17:54:32-0400

"\sum{x} \to 4.5-7.5\cdot \cos 30°+3\cdot \cos40°=0.303 m"


"\sum{y} \to 7.5\cdot \sin 30°-3\cdot \sin40°=1.82 m"


"Displacement =\sqrt{0.303^2+1.82^2}=1.85m"


"\overrightarrow{r}=0.303\overrightarrow{i}+1.82\overrightarrow{j}"